OpenStudy (anonymous):

Calculate the following integral \[\int\limits \tan^{-1} x dx \]

OpenStudy (mimi_x3):

integration by parts u = tan^(-1)x du / dx = 1/(1+x^2) dv = 1 v = x

OpenStudy (mimi_x3):

\[= xtan^{-1}x - \int x * \frac{1}{1+x^2}dx\]

OpenStudy (mimi_x3):

\[ = xtan^{-1}x - \int\frac{x}{1+x^2}dx\]

OpenStudy (mimi_x3):

for \[\int \frac{x}{1+x^2}\] use u-sub u = 1+ x^2 du/dx = 2x

OpenStudy (anonymous):

0.5ln1+x^2

OpenStudy (anonymous):

I've got the answer it says so thank you!